# Computing a Fixed Point of Contraction Maps in Polynomial Queries

### v1

@@ 2 Preliminaries
Let $10<\gamma<$ 1 and $(\mathcal{M},d)$ be a metric space. A map $f:\mathcal{M}\mapsto\mathcal{M}$ is a $(1-\gamma)$-contraction map with respect to $(\mathcal{M},d)$ if $d(f(x),f(y))\leq(1-\gamma)\cdot d(x,y)$ for all $x,y\in\mathcal{M}$.
@@
Every non-expansive map has a fixed point, i.e., $x^{*}$ with $f(x^{*})=x^{*}$, and it is unique when $f$ is a $(1-\gamma)$-contraction map for any $\gamma>0$. In this paper, we study the query complexity of finding an $\varepsilon$-fixed point of a $(1-\gamma)$-contraction map $f$ over the $k$-cube $[0,1]^{k}$ with respect to the infinity norm:
@@
In problem $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$, we are given oracle access to a $(1-\gamma)$-contraction map $f$ over $[0,1]^{k}$ with respect to the infinity norm, i.e., $f$ satisfies
@@
We define similar problems for non-expansive maps over $[0,1]^{k}$:
@@
We write $\textsc{StrNonExp}_{\infty}(\varepsilon,k)$ to denote the problem with the same input but the goal is to find a strong $\varepsilon$-fixed point of $f$.
@@
Notation. Given a positive integer $m$, we use $[m]$ to denote $\{1,\ldots,m\}$. For a real number $t\in\mathbb{R}$, we let ${\rm sgn}(t)=1$ if $t>0$, ${\rm sgn}(t)=-1$ if $t<0$, and ${\rm sgn}(t)=0$ if $t=0$. Given positive integers $n$ and $k$, we use $\texttt{OE}(n,k)$ to denote the set of all integer points $x\in[0,n]^{k}$ such that either $x_{i}$ is odd for all $i\in[k]$ or $x_{i}$ is even for all $i\in[k]$. (OE is a shorthand for odd and even points.)
For a point $x\in\mathbb{R}^{k}$ (not necessarily in $[0,n]^{k}$), a coordinate $i\in[k]$, and a sign $\phi\in\left\{\pm 1\right\}$, we use $\mathcal{P}_{i}(x,\phi)$ to denote
@@
Given a $(1-\gamma)$-contraction map $g$ over $[0,n]^{k}$, we use $\texttt{Fix}(g)$ to denote the unique fixed point of $g$. For any point $x\in[0,n]^{k}$, we use $\texttt{Around}(x)$ to denote the set
$\texttt{Around}(x):=\big{\{}y\in\texttt{OE}(n,k):\|x-y\|_{\infty}\leq 1\big{\}}.$
@@
$\|g(y)-y\|_{\infty}\leq\|g(y)-g(x)\|_{\infty}+\|x-y\|_{\infty}\leq 2,\quad\text{for any $y\in\texttt{Around}(x)$}.$
@@ 3 Characterizing the Unique Fixed Point
Putting ()($\star$)( ⋆ ) and ()($\diamond$)( ⋄ ) together and the facts that $\|c-a\|_{\infty}=4$ and $c_{j}=a_{j}$, we have
@@
Since we are considering points not in $\smash{\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(c,s_{i})}$, it must be the case that $x\in\mathcal{P}_{j}(c,-s_{j})$ and thus, $x_{j}\leq c_{j}$. Since $\smash{\|x-a\|_{\infty}>{8}/{\gamma}}$, we have $x_{j}\leq a_{j}$; otherwise $a_{j}\leq x_{j}\leq c_{j}$ and thus,
@@
Next, we show an upper bound on $\|x-a\|_{\infty}$. Recall that $x\in\mathcal{P}_{j}(c,-s_{j})$. Consider $y=x-4s$. We have $y\in\mathcal{P}_{j}(a,-s_{j})$. So
@@
$s_{j}(y_{j}-b_{j})\geq y_{i^{*}}-b_{i^{*}}\geq 0.$
Our goal is to show $y\in\mathcal{P}_{j}(b,s_{j})$ and it suffices for us to show $|y_{j}-b_{j}|\geq|y_{i}-b_{i}|$ for all $i\in[k]$.
@@
Thus $x\notin\mathcal{P}_{j}(a,-1)$ and $x\notin\mathcal{P}_{j}(a,+1)$ This finishes the proof of the lemma. ∎
@@ 4 The Algorithm
1:Let $\texttt{Cand}^{0}\leftarrow\texttt{OE}(n,k)$
@@
3:     Find and query an $a^{t}\in\texttt{OE}(n,k)$ such that $a^{t}$ is a balanced point of $\texttt{Cand}^{t-1}$ on some $j\in[k]$:
[…]
@@
We prove Theorem 1 in this section. Our algorithm for $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is described in Algorithm 1. Given oracle access to a $(1-\gamma)$-contraction map $g:[0,n]^{k}\rightarrow[0,n]^{k}$, we show that it can find a $({16}/{\gamma})$-fixed point of $g$ within
$O\left(k^{2}\log\left(\frac{1}{\varepsilon\gamma}\right)\right)$
many queries. This is sufficient given Lemma 1.
The analysis of Algorithm 1 uses the following theorem which we prove in the next section. In particular, it guarantees the existence of the point $a^{t}$ to be queried in round $t$ that satisfies (3). We will call a point $q^{*}$ with the property stated in Algorithm 1 below a balanced point for $T$.
@@
theoremexistenceofbalancedpoint For any $T\subseteq\emph{{OE}}(n,k)$, there exist a point $q^{*}\in\emph{{OE}}(n,k)$ and an $i^{*}\in[k]$ such that
$\big{|}\mathcal{P}_{i^{*}}(q^{*},\phi)\cap T\big{|}\geq\frac{1}{2k}\cdot|T|,\quad\text{for both $\phi\in\left\{\pm 1\right\}$}.$
At a high level, Algorithm 1 maintains a subset of grid points $\texttt{OE}(n,k)$ as candidate solutions, which is denoted by $\texttt{Cand}^{t}$ after round $t$. We show the following invariants:
@@
[…]
@@
We start with the proof of (5). Suppose that $a^{t}$ satisfies (3) with $j\in[k]$. Then by Lemma 4, there must exist a sign $\phi\in\left\{\pm 1\right\}$ such that
$\mathcal{P}_{j}(a^{t},\phi)\cap\left(\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(b^{t},s_{i})\right)=\emptyset.$
The inequality (5) follows directly from (3).
Next we prove by induction that ${\texttt{Around}}({\texttt{Fix}}(g))\subseteq{\texttt{Cand}}^{t}$ for every $t$ before the round that the algorithm terminates. The basis is trivial given that $\texttt{Cand}^{0}$ is set to be $\texttt{OE}(n,k)$. For round $t\geq 1$, we assume that $\|g(a^{t})-a^{t}\|_{\infty}>{16}/{\gamma}$; otherwise a solution is found and the algorithm terminates.
@@
We finish the proof by using the inductive hypothesis $\texttt{Around}(\texttt{Fix}(g))\subseteq\texttt{Cand}^{t-1}$ and (4). ∎
Given Lemma 5 and that $|\texttt{Cand}^{0}|\leq n^{k}$, within at most
$O\left(k\log(n^{k})\right)=O\left(k^{2}\log\left(\frac{1}{\varepsilon\gamma}\right)\right)$
@@ 5 Existence of Balanced Point
We prove Algorithm 1 in this section. Before that, we first illustrate why the structure of $\texttt{OE}(n,k)$ is necessary to guarantee the existence of a balanced point. In particular, for the standard grid $[n]^{k}$, we give an example of a set $T\subseteq[n]^{k}$ such that no point in $[n]^{k}$ can be a balanced point.
Example 1.
Let $p=(n/2,\ldots,n/2)$ if $n$ is odd and let $p=(\frac{n+1}{2},\cdots,\frac{n+1}{2})$ if $n$ is even. Note that $p\notin[n]^{k}$. We construct the following $T\subseteq[n]^{k}$ such that $p$ is the only choice of a balanced point:
$T=\left\{x\in[n]^{k}:|x_{i}-p_{i}|=\|x-p\|_{\infty}\text{ for all }i\in[k]\right\}.$
Precisely, it is easy to verify that for any point $p^{\prime}\neq p$ and any coordinate $i\in[k]$, we have that
$\big{|}\mathcal{P}_{i^{*}}(p^{\prime},\phi)\cap T\big{|}\leq O\left(|T|/2^{k}\right),\quad\text{for some $\phi\in\left\{\pm 1\right\}$}.$
However, since $p\notin[n]^{k}$, we conclude that there is no balanced point for $T$ in $[n]^{k}$.
@@
For every integer $t\geq 4$, there exist $p^{*}\in[-1/4,n+1/4]^{k}$ and a coordinate $i^{*}\in[k]$ such that $\emph{{vol}}(\mathcal{P}_{i^{*}}(p^{*},+1)\cap S^{t})=\emph{{vol}}(\mathcal{P}_{i^{*}}(p^{*},-1)\cap S^{t})\geq\emph{{vol}}(S^{t})/2k$.
@@
$0\leq\frac{\textbf{vol}\big{(}\mathcal{P}_{i}(p,+1)\cap S^{t}\big{)}}{(n+0.5)^{k-1}}\leq n+\frac{1}{4}-p_{i}\quad\text{and}\quad 0\leq\frac{\textbf{vol}\big{(}\mathcal{P}_{i}(p,-1)\cap S^{t}\big{)}}{(n+0.5)^{k-1}}\leq p_{i}+\frac{1}{4}.$
@@
for all $i\in[k]$. On the other hand, we have
$\sum_{i\in[k]}\Big{(}\textbf{vol}\big{(}\mathcal{P}_{i}(p^{*},+1)\cap S^{t}\big{)}+\textbf{vol}\big{(}\mathcal{P}_{i}(p^{*},-1)\cap S^{t}\big{)}\Big{)}=\textbf{vol}(S^{t}).$
Therefore there must be an $i^{*}\in[k]$ such that
$\textbf{vol}\big{(}\mathcal{P}_{i^{*}}(p^{*},+1)\cap S^{t}\big{)}=\textbf{vol}\big{(}\mathcal{P}_{i^{*}}(p^{*},-1)\cap S^{t}\big{)}\geq\frac{\textbf{vol}(S^{t})}{2k}.$
This finishes the proof of the lemma. ∎
By Lemma 6, we have that for every $t$, there exist $p^{t}\in\left[-1/4,n+1/4\right]^{k}$ and $i_{t}\in[k]$ such that
$\textbf{vol}\big{(}\mathcal{P}_{i_{t}}(p^{t},+1)\cap S^{t}\big{)}=\textbf{vol}\big{(}\mathcal{P}_{i_{t}}(p^{t},-1)\cap S^{t}\big{)}\geq\frac{\textbf{vol}(S^{t})}{2k}.$
Given that there are only $k$ choices for $i_{t}$, there exists an $i^{*}\in[k]$ such that $\{p^{t}\}$ has an infinite subsequence $\{p^{t_{\ell}}\}_{\ell\geq 1}$ with $t_{1}<t_{2}<\cdots$ such that
$\textbf{vol}\big{(}\mathcal{P}_{i^{*}}(p^{t_{\ell}},+1)\cap S^{t_{\ell}}\big{)}=\textbf{vol}\big{(}\mathcal{P}_{i^{*}}(p^{t_{\ell}},-1)\cap S^{t_{\ell}}\big{)}\geq\frac{\textbf{vol}(S^{t_{\ell}})}{2k}$
for all $\ell\geq 1$. Given that $\left[-1/4,n+1/4\right]^{k}$ is compact, $\{p^{t_{\ell}}\}$ has an infinite subsequence that converges to a point $p^{*}\in\left[-1/4,n+1/4\right]^{k}$. For convenience, we still refer to the subsequence as $\{p^{t_{\ell}}\}_{\ell\geq 1}$.
In Lemma 7, we show that both $\mathcal{P}_{i^{*}}(p^{*},+1)\cap T$ and $\mathcal{P}_{i^{*}}(p^{*},-1)\cap T$ are at least $|T|/2k$. After this, in Lemma 8, we show how to round $p^{*}$ to $q^{*}\in\texttt{OE}(n,k)$ while making sure that
$\mathcal{P}_{i^{*}}(p^{*},+1)\cap T\subseteq\mathcal{P}_{i^{*}}(q^{*},+1)\cap T\quad\text{and}\quad\mathcal{P}_{i^{*}}(p^{*},-1)\cap T\subseteq\mathcal{P}_{i^{*}}(q^{*},-1)\cap T.$
@@
$\min\left\{\frac{|\mathcal{P}_{i^{*}}(p^{*},+1)\cap T|}{|T|},\frac{|\mathcal{P}_{i^{*}}(p^{*},-1)\cap T|}{|T|}\right\}\geq\frac{1}{2k}.$
@@
Consider the easier case when $A$ is empty, i.e. $p_{i}^{*}+p_{j}^{*}$ and $p_{i}^{*}-p_{j}^{*}$ are integers for all $i\neq j\in[k]$. Let $\ell$ be a sufficiently large integer such that $1/t_{\ell}\leq 0.1$ and $\|p^{t_{\ell}}-p^{*}\|_{\infty}\leq 0.1$. We show that
$\frac{|\mathcal{P}_{i^{*}}(p^{*},+1)\cap T|}{|T|}\geq\frac{\textbf{vol}(\mathcal{P}_{i^{*}}(p^{t_{\ell}},+1)\cap S^{t_{\ell}})}{\textbf{vol}(S^{t_{\ell}})}\geq\frac{1}{2k}$
and the same inequality holds for the $-1$ side. For this purpose it suffices to show that every point $x\in T$ satisfies that
$\mathcal{P}_{i^{*}}(p^{t_{\ell}},+1)\cap B(x,1/t_{\ell})\neq\emptyset\ \ \Longrightarrow\ \ x\in\mathcal{P}_{i^{*}}(p^{*},+1).$
@@
Let’s prove the contrapositive so take any $x\notin\mathcal{P}_{i^{*}}(p^{*},+1)$. There exists a $j\neq i^{*}$ such that either
$x_{i^{*}}-p^{*}_{i^{*}}<x_{j}-p^{*}_{j}\quad\text{or}\quad x_{i^{*}}-p^{*}_{i^{*}}<p^{*}_{j}-x_{j}$
For the first case we have $x_{i^{*}}-x_{j}<p^{*}_{i^{*}}-p^{*}_{j}$. Since both sides are integers we have
$x_{i^{*}}-x_{j}\leq p^{*}_{i^{*}}-p^{*}_{j}-1$ (6)

### v2

@@ 2 Preliminaries
Let $0<\gamma<1$ and $(\mathcal{M},d)$ be a metric space. A map $f:\mathcal{M}\mapsto\mathcal{M}$ is a $(1-\gamma)$-contraction map with respect to $(\mathcal{M},d)$ if $d(f(x),f(y))\leq(1-\gamma)\cdot d(x,y)$ for all $x,y\in\mathcal{M}$.
@@
Every contraction map has a unique fixed point, i.e., $x^{*}$ with $f(x^{*})=x^{*}$ by Banach’s theorem. Every non-expansive map over $0,1]^{k}$ with respect to the infinite norm has a (not necessarily unique) fixed point by Brouwer’s theorem. In this paper, we study the query complexity of finding an $\varepsilon$-fixed point of a $(1-\gamma)$-contraction map $f$ over the $k$-cube $0,1]^{k}$ with respect to the infinity norm:
@@
In problem $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$, we are given oracle access to a $(1-\gamma)$-contraction map $f$ over $0,1]^{k}$ with respect to the infinity norm, i.e., $f$ satisfies
@@
We define similar problems for non-expansive maps over $0,1]^{k}$:
@@
We write $\textsc{StrNonExp}_{\infty}(\varepsilon,k)$ to denote the problem with the same input but the goal is to find a strong $\varepsilon$-fixed point of $f$, namely, a point $x$ such that $\|x-x^{*}\|_{\infty}\leq\varepsilon$, where $x^{*}$ is any fixed point of $f$.
@@
Notation. Given a positive integer $m$, we use $[m]$ to denote $\{1,\ldots,m\}$ and $[0:m]$ to denote $\left\{0,\ldots,m\right\}$. For a real number $t\in\mathbb{R}$, we let ${\rm sgn}(t)=1$ if $t>0$, ${\rm sgn}(t)=-1$ if $t<0$, and ${\rm sgn}(t)=0$ if $t=0$. Given positive integers $n$ and $k$, we use $\texttt{EVEN}(n,k)$ to denote the set of all integer points $x\in[0,n]^{k}$ such that $x_{i}$ is even for all $i\in[k]$.
For a point $x\in\mathbb{R}^{k}$ (not necessarily in $0,n]^{k}$), a coordinate $i\in[k]$, and a sign $\phi\in\left\{\pm 1\right\}$, we use $\mathcal{P}_{i}(x,\phi)$ to denote
@@
Given a $(1-\gamma)$-contraction map $g$ over $0,n]^{k}$, we use $\texttt{Fix}(g)$ to denote the unique fixed point of $g$. For any point $x\in[0,n]^{k}$, we use $\texttt{Around}(x)$ to denote the set
$\texttt{Around}(x):=\big{\{}y\in[0,n]^{k}:\|x-y\|_{\infty}\leq 1\big{\}}.$
@@
$\|g(y)-y\|_{\infty}\leq\|g(y)-g(x)\|_{\infty}+\|x-y\|_{\infty}\leq(2-\gamma)\|x-y\|_{\infty}\leq 2,\quad\text{for any $y\in\texttt{Around}(x)$}.$
Note that for any $x\in[0,n]^{k}$, $\texttt{Around}(x)\cap\texttt{EVEN}(n,k)$ is non-empty.
@@ 3 Characterizing the Unique Fixed Point
In this section, we prove three geometrical lemmas for characterizing the exact fixed point $\texttt{Fix}(g)$. The corresponding geometrical illustrations are in Figure 2, Figure 3, and Figure 4.
@@
aag(a)g(a)$a+4s$
(a) $a$ and $g(a)$ satisfy $s=(+1,+1)$.
aag(a)g(a)$a+4s$
(b) $a$ and $g(a)$ satisfy $s=(0,+1)$.
Figure 2: An illustration of Lemma 2. The light blue region is given by $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(a+4s,s_{i})$, which contains $\texttt{Fix}(g)$ (the unique fixed point of $g$).
@@
Putting $(\star)$ and $(\diamond)$ together and the facts that $\|c-a\|_{\infty}=4$ and $c_{j}=a_{j}$, we have
@@
Since we are considering points not in $\smash{\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(c,s_{i})}$, it must be the case that $x\in\mathcal{P}_{j}(c,-s_{j})$ and thus, $x_{j}\leq c_{j}$. Since $\smash{\|x-a\|_{\infty}>{8}/{\gamma}}$, we have $x_{j}\leq a_{j}$; otherwise $a_{j}\leq x_{j}\leq c_{j}$ and $\|x-c\|_{\infty}\leq c_{j}-x_{j}\leq c_{j}-a_{j}\leq 4$ and thus,
@@
Next, we show an upper bound on $\|x-a\|_{\infty}$. Recall that $x\in\mathcal{P}_{j}(c,-s_{j})$. Consider y=x4sy=x-4s. We have $y\in\mathcal{P}_{j}(a,-s_{j})$. So
@@
aabbb+2sb+2sxx
(a) The case $s=(+1,+1)$.
aabbb+2sb+2sxx
(b) The case $s=(0,+1)$.
Figure 3: An illustration of Lemma 3. The light blue region is given by $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(b+2s,s_{i})$, which contains $x$. $\texttt{Around}(x)$ is the small square around $x$, which is contained in $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(b,s_{i})$.
@@
$s_{j}(y_{j}-b_{j})\geq y_{i^{*}}-b_{i^{*}}\geq 0,$
thus $s_{j}(y_{j}-b_{j})=|y_{j}-b_{j}|$. Our goal is to show $y\in\mathcal{P}_{j}(b,s_{j})$ and it suffices for us to show $|y_{j}-b_{j}|\geq|y_{i}-b_{i}|$ for all $i\in[k]$.
@@
aaa+2sa+2s
(a) The case $s=(+1,+1)$. The light blue region is given by $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(a+2s,s_{i})$. Clearly it does not intersect with $\mathcal{P}_{1}(a,-1)$ or $\mathcal{P}_{2}(a,-1)$.
aaa+2sa+2s
(b) The case $s=(0,+1)$. The light blue region is given by $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(a+2s,s_{i})$. Clearly it does not intersect with $\mathcal{P}_{2}(a,-1)$, $\mathcal{P}_{1}(a,-1)$, or $\mathcal{P}_{1}(a,1)$.
Figure 4: An illustration of Lemma 4. When $s_{i}\neq 0$, we can prove that $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(a+2s,s_{i})$ does not intersect with $\mathcal{P}_{i}(a,-s_{i})$; see left figure above. When $s_{i}=0$, we can prove $\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(a+2s,s_{i})$ does not intersect with $\mathcal{P}_{i}(a,-1)$ or $\mathcal{P}_{i}(a,1)$; see right figure above.
@@
Thus $x\notin\mathcal{P}_{j}(a,-1)$ and $x\notin\mathcal{P}_{j}(a,+1)$. This finishes the proof of the lemma. ∎
@@ 4 The Algorithm
1:Let $\texttt{Cand}^{0}\leftarrow\texttt{EVEN}(n,k)$.
@@
3:  Find and query an $a^{t}\in[0:n]^{k}$ such that $a^{t}$ is a balanced point of $\texttt{Cand}^{t-1}$.
@@
We prove Theorem 1 in this section. Our algorithm for $\textsc{Contraction}_{\infty}(\varepsilon,\gamma,k)$ is described in Algorithm 1. Given oracle access to a $(1-\gamma)$-contraction map $g:[0,n]^{k}\rightarrow[0,n]^{k}$, we show that it can find a $({16}/{\gamma})$-fixed point of $g$ within $O\left(k\log\left(n\right)\right)$ many queries. This is sufficient given Lemma 1.
The analysis of Algorithm 1 uses the following theorem which we prove in the next section. In particular, it guarantees the existence of the point $a^{t}$ to be queried in round $t$ that satisfies that for any $s\in\left\{\pm 1\right\}^{k}$, we have $\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(a^{t},s_{i})\right)$ intersects with at least half of $\texttt{Cand}^{t-1}$. We will call a point $q^{*}$ with the property stated in Algorithm 1 below a balanced point for $T$.
@@
theoremexistenceofbalancedpoint For any $T\subseteq\emph{{EVEN}}(n,k)$, there exist a point $q^{*}\in[0:n]^{k}$ such that
$\left|T\cap\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(q^{*},s_{i})\right)\right|\geq\frac{1}{2}\cdot|T|,\quad\text{for all $s\in\left\{\pm 1\right\}^{k}$}.$
At a high level, Algorithm 1 maintains a subset of grid points $\texttt{EVEN}(n,k)$ as candidate solutions, which is denoted by $\texttt{Cand}^{t}$ after round $t$. We show the following invariants:
@@
For every round $t\geq 1$, either the point $a^{t}$ queried is a $(16/\gamma)$-fixed point of $g$ (and the algorithm terminates), or we have both $(\emph{{Around}}(\emph{{Fix}}(g))\cap\emph{{EVEN}}(n,k))\subseteq\emph{{Cand}}^{t}$ and
$\big{|}\emph{{Cand}}^{t}\big{|}\leq\frac{1}{2}\cdot\big{|}\emph{{Cand}}^{t-1}\big{|}.$ (3)
@@
We start with the proof of (3). Suppose that $a^{t}$ is a balanced point of $\texttt{Cand}^{t-1}$. Then by Lemma 4, for any coordinate $j\in[k]$, there must exist a sign $\phi_{j}\in\left\{\pm 1\right\}$ such that
$\mathcal{P}_{j}(a^{t},\phi_{j})\cap\left(\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(b^{t},s_{i})\right)=\emptyset.$
Since $\texttt{Cand}^{t}\coloneqq\texttt{Cand}^{t-1}\cap\left(\bigcup_{i:s_{i}\neq 0}\mathcal{P}_{i}(b^{t},s_{i})\right)$, we know that for every coordinate $j\in[k]$, there exists $\phi_{j}\in\left\{\pm 1\right\}$ such that $\texttt{Cand}^{t}\cap\mathcal{P}_{j}(a^{t},\phi_{j})=\emptyset$. The inequality (3) follows directly from the definition of the balanced point.
Next we prove by induction that $({\texttt{Around}}({\texttt{Fix}}(g))\cap\texttt{EVEN}(n,k))\subseteq{\texttt{Cand}}^{t}$ for every $t$ before the round that the algorithm terminates. The basis is trivial given that $\texttt{Cand}^{0}$ is set to be $\texttt{EVEN}(n,k)$. For round $t\geq 1$, we assume that $\|g(a^{t})-a^{t}\|_{\infty}>{16}/{\gamma}$; otherwise a solution is found and the algorithm terminates.
@@
We finish the proof by using the inductive hypothesis $(\texttt{Around}(\texttt{Fix}(g))\cap\texttt{EVEN}(n,k))\subseteq\texttt{Cand}^{t-1}$ and (2). ∎
Recall that ${\texttt{Around}}({\texttt{Fix}}(g))\cap\texttt{EVEN}(n,k)$ is non-empty. Given Lemma 5 and that $|\texttt{Cand}^{0}|\leq n^{k}$, within at most
$O\left(\log(n^{k})\right)=O\left(k\log\left(\frac{1}{\varepsilon\gamma}\right)\right)$
@@ 5 Existence of Balanced Point
For every integer $t\geq 4$, there exist $p^{*}\in[-1/4,n+1/4]^{k}$ such that $\emph{{vol}}(\mathcal{P}_{i}(p^{*},+1)\cap S^{t})=\emph{{vol}}(\mathcal{P}_{i}(p^{*},-1)\cap S^{t})$ holds for all coordinates $i\in[k]$.
@@
$0\leq\frac{\textbf{vol}\big{(}\mathcal{P}_{i}(p,+1)\cap S^{t}\big{)}}{(n+0.5)^{k-1}}\leq n+\frac{1}{4}-p_{i}\quad\text{and}\quad 0\leq\frac{\textbf{vol}\big{(}\mathcal{P}_{i}(p,-1)\cap S^{t}\big{)}}{(n+0.5)^{k-1}}\leq p_{i}+\frac{1}{4},$
where the upper bounds above simply follow from the fact that
$\textbf{vol}\big{(}\mathcal{P}_{i}(p,+1)\big{)}\leq\bigg{(}n+\frac{1}{4}-p_{i}\bigg{)}\cdot(n+0.5)^{k-1}\quad\text{and}\quad\textbf{vol}\big{(}\mathcal{P}_{i}(p,-1)\big{)}\leq\bigg{(}p_{i}+\frac{1}{4}\bigg{)}\cdot(n+0.5)^{k-1}.$
@@
for all $i\in[k]$. This finishes the proof of the lemma. ∎
By Lemma 6, we have that for every $t$, there exists $p^{t}\in\left[-1/4,n+1/4\right]^{k}$ such that
$\textbf{vol}\big{(}\mathcal{P}_{i}(p^{t},+1)\cap S^{t}\big{)}=\textbf{vol}\big{(}\mathcal{P}_{i}(p^{t},-1)\cap S^{t}\big{)}$
for all $i\in[k]$. Given that $\left[-1/4,n+1/4\right]^{k}$ is compact, $\{p^{t}\}_{t\geq 1}$ has an infinite subsequence that converges to a point $p^{*}\in\left[-1/4,n+1/4\right]^{k}$. We refer to the subsequence as $\{p^{t_{\ell}}\}_{\ell\geq 1}$.
In Lemma 7, we show that $p^{*}$ is already a desired balanced point but it may be off-grid. In Lemma 8, we show how to round $p^{*}$ to $q^{*}\in[0:n]^{k}$ while making sure that
$\mathcal{P}_{i}(p^{*},+1)\cap T\subseteq\mathcal{P}_{i^{*}}(q^{*},+1)\cap T\quad\text{and}\quad\mathcal{P}_{i^{*}}(p^{*},-1)\cap T\subseteq\mathcal{P}_{i^{*}}(q^{*},-1)\cap T.$
@@
$\left|\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{*},s_{i})\right)\cap T\right|\geq\frac{1}{2}\cdot|T|,\quad\text{for all $s\in\left\{\pm 1\right\}^{k}$}.$
@@
Consider the easier case when $A$ is empty, i.e. $p_{i}^{*}+p_{j}^{*}$ and $p_{i}^{*}-p_{j}^{*}$ are integers for all $i\neq j\in[k]$. Let $\ell$ be a sufficiently large integer such that $1/t_{\ell}\leq 0.1$ and $\|p^{t_{\ell}}-p^{*}\|_{\infty}\leq 0.1$. Fixing any $s\in\left\{\pm 1\right\}^{k}$, we first show the following claim:
Claim 1.
For every point $x\in T$, we have
$\left(\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\right)\cap B(x,1/t_{\ell})\neq\emptyset\ \ \Longrightarrow\ \ x\in\bigcup_{i\in[k]}\mathcal{P}_{i}(p^{*},s_{i}).$
Proof.
We show a slightly stronger statement in the proof: For any $i\in[k]$, we have
$\mathcal{P}_{i}(p^{t_{\ell}},s_{i})\cap B(x,1/t_{\ell})\neq\emptyset\ \ \Longrightarrow\ \ x\in\mathcal{P}_{i}(p^{*},s_{i}).$
@@
Let’s prove the contrapositive so take any $x\in T$ such that $x\notin\mathcal{P}_{i}(p^{*},s_{i})$. Assume without loss of generality that $s_{i}=+1$. We know that there exists a $j\neq i$ such that either
$x_{i}-p^{*}_{i}<x_{j}-p^{*}_{j}\quad\text{or}\quad x_{i}-p^{*}_{i}<p^{*}_{j}-x_{j}.$
For the first case we have $x_{i}-x_{j}<p^{*}_{i}-p^{*}_{j}$. Since both sides are integers we have
$x_{i}-x_{j}\leq p^{*}_{i}-p^{*}_{j}-1$ (4)